### v3

# Towards Unified Attribution in Explainable AI, Data-Centric AI, and Mechanistic Interpretability

Abstract
The increasing complexity of AI systems has made understanding their behavior critical. Numerous interpretability methods have been developed to attribute model behavior to three key aspects: input features, training data, and internal model components, which emerged from explainable AI, data-centric AI, and mechanistic interpretability, respectively. However, these attribution methods are studied and applied rather independently, resulting in a fragmented landscape of methods and terminology. This position paper argues that feature, data, and component attribution methods share fundamental similarities, and a unified view of them benefits both interpretability and broader AI research. To this end, we first analyze popular methods for these three types of attributions and present a unified view demonstrating that these seemingly distinct methods employ similar techniques (such as perturbations, gradients, and linear approximations) over different aspects and thus differ primarily in their perspectives rather than techniques. Then, we demonstrate how this unified view enhances understanding of existing attribution methods, highlights shared concepts and evaluation criteria among these methods, and leads to new research directions both in interpretability research, by addressing common challenges and facilitating cross-attribution innovation, and in AI more broadly, with applications in model editing, steering, and regulation.
1 Introduction
The escalating complexity of AI systems necessitates a deeper understanding of their behavior, giving rise to the field of AI interpretability. A central question in interpretability research is how to attribute an AI system’s behavior to its underlying factors. To address this question, researchers across different domains have developed various attribution methods. Feature attribution (FA), which emerged from explainable AI, quantifies the influence of input features on a model’s output at test time to reveal critical decision drivers (Zeiler and Fergus, 2014; Ribeiro et al., 2016; Horel and Giesecke, 2020, 2022; Lundberg and Lee, 2017; Smilkov et al., 2017). Data attribution (DA), a core approach in data-centric AI, analyzes how training data points shape overall model behavior during the training process (Koh and Liang, 2017; Ghorbani and Zou, 2019; Ilyas et al., 2022). Component attribution (CA), a key branch of mechanistic interpretability, examines how internal model components, such as neurons or layers within a neural network (NN), contribute to model behavior (Vig et al., 2020; Meng et al., 2022; Nanda, 2023; Shah et al., 2024). Although numerous attribution methods and comprehensive survey papers have been published (Guidotti et al., 2018; Covert et al., 2021; Wang et al., 2024; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024; Rai et al., 2024; Räuker et al., 2023; Vilas et al., 2024), these methods have largely been studied and applied in isolation by their respective communities, resulting in a fragmented landscape of methods and terminology for similar concepts (Saphra and Wiegreffe, 2024).
We take the position that 1) feature, data, and component attribution share core techniques despite their different perspectives, and that 2) this more unified view advances both interpretability, by overcoming common challenges and stimulating cross-attribution innovation, and broader AI research, with applications in model editing, steering, and regulation. In the following sections, we first define the attribution problem and formalize a unified framework that encompasses all three types of attributions (§2). Next, we examine each attribution type, analyzing its popular methods, and demonstrate how all three attribution types share fundamental techniques (i.e., perturbations, gradients, and linear approximations), key concepts, and evaluation criteria (§3). We summarize representative attribution methods based on their techniques in Table 1. Then, we demonstrate how this unified view can advance both interpretability and broader AI research (§4). We also discuss alternative views of the paper’s position (§5). In summary, we believe that this unified view enhances our understanding of attribution methods, bridges the current fragmented landscape, and provides new insights and research directions.
2 The attribution problem
Various attribution methods have been developed to analyze model behavior from different aspects. In this section, we formally introduce three types of attribution problems and show how they fall under a unified framework. Consider a learning problem with -dimensional input features . During training, a dataset of data points: is used to train a model with parameters and components by optimizing the loss function . At test (inference) time, the model generates an output for a new input . For notational simplicity, we omit and “test” and use and when the context is unambiguous. A notation summary is provided in Appendix A. The core objective of all three problems is to attribute the model’s output to different aspects and quantify their influence with attribution scores.
Feature attribution (FA) quantifies how input features affect the model’s output. A features may be pixels in images, tokens in text, or other domain-specific units with as its attribution score.
Data attribution (DA) analyzes how training data shape model behavior during training. We quantify the influence of each training point through its attribution score .
Component attribution (CA) studies the role of model components in producing model outputs. The component can have various definitions, like a neuron in an NN, with as its attribution score.
As illustrated in Figure 1, these three attribution problems share a fundamental connection: they all seek an attribution function that assigns scores to specific aspects (features , training points , or components ) for a given test output , differing only in the choice of aspects.
| Technique | Technique Details | Feature Attribution | Data Attribution | Component Attribution |
| Perturbation | Direct | Occlusions (Zeiler and Fergus, 2014) | LOO (Cook and Weisberg, 1982) | Causal Tracing (Meng et al., 2022) |
| RISE (Petsiuk, 2018) | Path Patching (Wang et al., 2022) | |||
| Vig et al. (2020) | ||||
| Bau et al. (2020) | ||||
| ACDC (Conmy et al., 2023) | ||||
| Game-Theoretic | SHAP (Lundberg and Lee, 2017) | Data Shapley (Ghorbani and Zou, 2019) | Neuron Shapley (Ghorbani and Zou, 2020) | |
| (Shapley) | TMC Shapley (Ghorbani and Zou, 2019) | |||
| KNN Shapley (Jia et al., 2019) | ||||
| Beta Shapley (Kwon and Zou, 2022) | ||||
| Game-Theoretic | STII (Dhamdhere et al., 2019) | Data Banzhaf (Wang and Jia, 2023) | – | |
| (Others) | BII (Patel et al., 2021) | |||
| Core Value (Yan and Procaccia, 2021) | ||||
| Myerson Value (Chen et al., 2018b) | ||||
| HN Value (Zhang et al., 2022) | ||||
| Mask Learning | Dabkowski and Gal (2017) | – | Csordás et al. (2020) | |
| L2X (Chen et al., 2018a) | Subnetwork Pruning (Cao et al., 2021) | |||
| Gradient | First-Order | Vanilla Gradients (Simonyan et al., 2013) | GradDot/GradCos (Pruthi et al., 2020) | Attribution Patching (Nanda, 2023) |
| Gradient Input (Shrikumar et al., 2017) | EAP (Syed et al., 2023) | |||
| SmoothGrad (Smilkov et al., 2017) | ||||
| GBP (Springenberg et al., 2014) | ||||
| Grad-CAM (Selvaraju et al., 2016) | ||||
| Second-Order | Integrated Hessian (Janizek et al., 2021) | IF (Koh and Liang, 2017) | – | |
| (Hessian/IF) | FastIF (Guo et al., 2021) | |||
| Arnoldi IF (Schioppa et al., 2022) | ||||
| EK-FAC (Grosse et al., 2023) | ||||
| RelateIF (Barshan et al., 2020) | ||||
| Tracing Path | Integrated Grad (Sundararajan et al., 2017) | TracIn (Pruthi et al., 2020) | Attribution Path Patching (Nanda, 2023) | |
| SGD-Influence (Hara et al., 2019) | ||||
| SOURCE (Bae et al., 2024) | ||||
| Linear Approximation | LIME (Ribeiro et al., 2016) | Datamodels (Ilyas et al., 2022) | COAR (Shah et al., 2024) | |
| C-LIME (Agarwal et al., 2021) | TRAK (Park et al., 2023) |
3 Unifying feature, data, and component attribution
3.1 Understanding feature attribution (FA)
FA quantifies how individual features of an input influence the model’s output through attribution scores . Applied to model inference at test time, it explains model behavior without altering model parameters. The attribution results can help with feature selection, model debugging, and improve user trust. FA methods can be broadly classified into three categories: perturbation-based, gradient-based, and linear approximation. We discuss some prominent methods in each category below and provide more details in Appendix C.
Perturbation-Based FA. These methods attribute feature importance by measuring how model outputs change when input features are modified and especially removed (also called removal-based methods (Covert et al., 2021)). Direct perturbation straightforwardly applies this idea, e.g., the pioneering Occlusion method (Zeiler and Fergus, 2014) in computer vision (CV) replaces image pixels with grey squares and measures prediction changes. For images, pixel attribution scores create a saliency map highlighting the most influential regions. RISE (Petsiuk, 2018) advances this method by perturbing multiple image regions and combining the results. While intuitive, direct perturbation fails to capture synergistic interactions between multiple features. Game-theoretic perturbation addresses this limitation by modeling features as game players collaborating toward the model’s output. The Shapley value (Shapley, 1953) provides a foundational solution within this framework and has inspired numerous FA methods (Sundararajan and Najmi, 2020). Computing Shapley value attributions involves measuring a specific type of perturbation: how adding a feature to different feature subsets changes the model’s output compared to the subset alone, known as the marginal contribution of to the subset. The final attribution score captures feature interactions by aggregating these marginal contributions across all possible feature subsets, which can be computationally challenging. Various approximation methods have been proposed for this computation, and KernelSHAP Lundberg and Lee (2017) (or simply SHAP) has gained widespread adoption because of its efficient kernel-based approximation. Since removing input features can be viewed as applying binary masks for features, perturbation mask learning methods use learnable masks representing feature removal probabilities, which offer more nuanced control compared to binary masks. A new mask can be learned for each input (Fong and Vedaldi, 2017) or only once as a global masking model (Dabkowski and Gal, 2017). The masking model acts as the attribution function . While initial training is required, can be faster at test time by generating masks through a single forward pass.
Gradient-based FA. Gradients of model outputs with respect to input features , , quantify output sensitivity to small input changes (Erhan et al., 2009; Baehrens et al., 2010), measuring feature influence without requiring perturbations. Gradient-based methods are more computationally efficient, needing only a single or a few forward and backward pass(es) to compute instead of at least model perturbations for features. Also originating from CV, these methods gained widespread adoption for generating image saliency maps (Simonyan et al., 2013) (also called sensitivity maps (Smilkov et al., 2017)). The “vanilla gradients” method uses the gradients of the output class (log)probability with respect to input pixels as attribution scores (Simonyan et al., 2013). Since then, numerous enhanced gradient-based methods has been proposed. For example, Integrated Gradients (Sundararajan et al., 2017) accumulates gradients along a path from a baseline to the actual input, and Integrated Hessians (Janizek et al., 2021) further extends the analysis to feature interactions by computing the Hessian matrix. SmoothGrad (Smilkov et al., 2017) generates multiple copies of the input with added Gaussian noise and computes sensitivity maps for each noisy sample. It then averages these maps to reduce noise while preserving salient features for model outputs.
Linear approximation for FA. These methods offer an alternative approach to FA by approximating the complex behavior of around an input of interest using a linear model, typically in the form of with coefficients and bias , and using coefficient as the FA score of feature . LIME (Ribeiro et al., 2016) exemplifies this approach. It fits a sparse linear model to capture local model behavior and uses binary indicators (0 or 1) rather than actual feature values as inputs to the linear model, only representing feature exclusion or inclusion. The resulting linear model coefficients explain how each feature’s presence influences the approximated model’s output. Notably, LIME can also be viewed through a perturbation lens, as it fundamentally perturbs input features to approximate the output. This connection suggests that different FA methods can be understood within a common mathematical framework, which we explore next.
3.1.1 Unifying feature attribution methods via local function approximation (LFA)
While FA methods can be grouped in the three categories above, many of them can be unified under the LFA framework (Han et al., 2022). In this framework, a model is approximated around a point of interest in a local neighborhood by an interpretable model using a loss function . Eight prominent FA methods (Occlusion, KernelSHAP, Vanilla Gradients, Gradients Input, Integrated Gradients, SmoothGrad, LIME, and C-LIME) are instances of this framework, differing only in their choices of local neighborhoods and loss functions (Appendix Table 3) Han et al. (2022).
The LFA framework (Han et al., 2022) enhances our understanding of FA methods in several ways. First, it provides conceptual coherence to the field. While different methods appear to have distinct motivations, this framework reveals their shared goal of LFA. Second, placing diverse methods under a single framework enables direct comparisons among them. This comparative lens allows us to better understand their similarities, differences, and behavior, such as why different methods can generate disagreeing explanations for the same model prediction (Krishna* et al., 2024). Third, this unification enables theoretical simplicity. Instead of studying methods individually, theoretical analyses can be performed using the framework and applied to each method Han et al. (2022); Bilodeau et al. (2024); Bordt and von Luxburg (2023). Fourth, the conceptual understanding brought about by unification leads to principled, practical recommendations (Han et al., 2022). Additional details on this unification are discussed in Appendix C.6.
3.2 Understanding data attributions (DA)
DA studies how the training dataset shapes model behavior during training. It is also known as data valuation, as it assesses the value of data from vendors. For each training data point , an attribution score traces back to training to quantify ’s influence on the model’s output for a test point . DA scores can be used to characterize data properties, help identify mislabeled data, and justify training data values. Like FA, DA uses the techniques of perturbations, gradients, and linear approximation. We examine prominent methods below, with additional details in Appendix D.
Perturbation-based DA. These DA methods examine changes in the model behavior after removing or re-weighting the training data points and subsequently retraining the model (as a result, they are also called retraining-based methods (Hammoudeh and Lowd, 2024)). Leave-One-Out (LOO) attribution, analogous to direct perturbation in FA, is a prominent example that trains a model on the complete training set and then separately removes each individual data point and retrains the model. The attribution score for each removed point is determined by the corresponding performance difference. The LOO approach has a long history in statistics (Cook and Weisberg, 1982) and has proven useful for DA for modern AI models (Jia et al., 2021). It provides valuable insights with its main limitation being the computational cost of model retraining for each data point. Many DA methods later can be viewed as efficient approximations of LOO. A natural extension of LOO is to leave a set of data points out to evaluate their collective impact (Ilyas et al., 2022). As the direct perturbation for FA, LOO overlooks interactions between data points, potentially missing subtle influence behaviors (Lin et al., 2022; Jia et al., 2021). This is addressed by game-theoretic DA methods like Data Shapley (Ghorbani and Zou, 2019) which applies the Shapley value to DA by computing each training point’s aggregated marginal contribution across all training subsets. Similar to game-theoretic FA, they face prohibitive computational costs, as each marginal contribution requires retraining and there are possible subsets. Many approximations have been proposed to address this challenge (Section D.2).
Gradient-based DA. These DA methods use the gradients of the loss with respect to both training data and test data to assess the impact of training point on model output . The GradDot and GradCos methods use the dot product and cosine similarity of these two gradients as attribution scores Charpiat et al. (2019). Influence Function (IF), a classic statistical method for analyzing influential points in linear regression (Cook and Weisberg, 1980), has been adapted for modern AI models (Koh and Liang, 2017). IF approximates LOO model parameter changes by Taylor expansion, computing both the gradient and the (inverse) Hessian, thus avoiding retraining. It offers an effective and computationally feasible alternative to LOO but also faces challenges including convexity assumptions and Hessian matrix computation which many follow-up methods try to address (Section D.3). Aside from these IF-based methods, TracIn (Pruthi et al., 2020) attributes the data influence throughout the entire training path by computing dot products between training and test data gradients at each training step from the initial model parameters to the final model parameters, accumulating these to capture a total influence across the training path. This path tracing method avoids limitations of LOO and IF, such as assigning identical attribution scores to duplicate training data points.
Linear approximation for DA. Datamodel (Ilyas et al., 2022) applies linear approximation to DA, similar to LIME in FA. It constructs a linear model with coefficients and vectors as inputs, where each input represents a subset of training data. is learned to map any subset of training data to output , where is trained on this subset with the given model architecture and training algorithm. The coefficients of thus represent the attribution scores of the data points. While Datamodel can effectively capture model behavior, constructing this large linear model with coefficients requires extensive counterfactual data obtained by training model on various subsets, making it computationally intensive. TRAK (Park et al., 2023) addresses the computational challenge by estimating Datamodels in a transformed space where the learning problem becomes convex. Furthermore, both approaches can be viewed as perturbation-based methods, similar to LIME, as they vary training data to construct linear models.
3.3 Understanding component attributions (CA)
CA studies the role of internal model components in model outputs. It plays a central role in the emerging mechanistic interpretability research which seeks to understand AI models by reverse engineering their internal mechanisms into interpretable algorithms. CA operates primarily at test time for model inference, quantifying how each model component contributes to the output through an attribution score . Components can be defined flexibly - from individual neurons and attention heads to entire layers and circuits (subnetworks). Like FA and DA, CA leverages the same three techniques: perturbations, gradients, and linear approximation. We examine key CA methods below and provide additional details in Appendix E.
Perturbation-based CA. Perturbation-based methods for CA are fundamentally similar to those for FA and DA. Components of the model, whether neurons, layers, or circuits, are perturbed to measure their effect on model behavior. Generally, the perturbations are chosen carefully to attempt to localize behaviors related to specific tasks or concepts. Causal mediation analysis (Pearl, 2022; Vig et al., 2020) is a perturbation-based method that leverages the abstraction of models to causal graphs, with graphs nodes corresponding to components, and graph edges represent the causal relationships between nodes. Causal mediation analysis is defined by an input cause and an output effect that is mediated by intermediate causal nodes, , between and . By perturbing , changes in can be measured to get attribution scores . Specifically, the effects of perturbations on component activations are measured during three separate runs: a clean run with no perturbations, a corrupted run where components are perturbed (may also be perturbing inputs to induce corrupted component activations), and a corrupted-with-restoration run that measures how much restoring a single component activation of the corrupted run by its original clean-run value can restore the output. The corrupted run can be repeated multiple times with different random noise added to obtain more robust CA scores. This analysis is frequently referred to as causal tracing (Meng et al., 2022) or activation patching, and path patching (Wang et al., 2022) when patching is applied to paths connecting components. Game-theoretic CA (Ghorbani and Zou, 2020) follows a similar approach to game-theoretic FA and DA. In particular, Neuron Shapley (Ghorbani and Zou, 2020) extends prior works on Shapley values to CA and mitigates computational cost through sampling and a multi-armed bandit algorithm to identifies neurons with large attribution scores. Mask Learning and Subnetwork Probing (Csordás et al., 2020; Cao et al., 2021) adopts a similar concept to FA mask learning, attempting to approximate either the model’s or a probe’s performance on a given task by searching for components that equivalently perform that task. More specifically, subnetwork probing optimizes a mask for the model weights as a continuous relaxation for searching, which essentially prunes the model.
Gradient-based CA. To improve the computational efficiency of CA methods, alterations of causal tracing leverage gradient-based approximations to do only two forward passes and a single backward pass to generate attributions. Attribution patching (Nanda, 2023) approximates the local change when patching a single activation from the corrupted to clean input. This is achieved by computing the backward pass for the corrupted output with respect to the patching metric and storing the gradients with respect to the activations. Finally, the difference between the clean and corrupted activations is taken and multiplied by the cached gradients to obtain attribution scores.
Linear approximation for CA. CA also employs linear approximations like LIME in FA and Datamodels in DA. COAR (Shah et al., 2024) attempts to decompose model behavior in terms of various components by predicting the counterfactual impact of ablating each component. Given the computational complexity, COAR employs linear approximations by assigning scores to each component and estimating the counterfactual effect of removing sets of components by summing their scores. Thus, the complexity of relationships between components is abstracted away.
3.4 Common concepts among attribution methods
As demonstrated in the previous sections, FA, DA, and CA methods leverage three shared techniques: perturbations, gradients, and linear approximations. Additional details are discussed in Appendix B and a summary of the methods is provided in Table 1. Beyond shared techniques, conceptual ideas also transfer across the three types of attributions. One such idea is the deliberate introduction of randomness and smoothing to enhance attribution robustness. This idea has proven effective across all three attribution types, with SmoothGrad using random noisy samples to denoise gradients for FA, TRAK ensembling attributions over multiple retrainings for DA, and causal mediation analysis aggregating results from multiple corrupted runs for CA. Another shared conceptual idea is tracking and aggregating results along paths. For example, Integrated Gradients calculates the path integral from base input to target input in FA, TracIn traces training paths to reveal dynamic data influences in DA, and path patching tracks component effects along residual stream paths in CA. These shared concepts are high-level ideas, not specific algorithms, and the definitions of paths can be different, but they highlight fundamental connections among different types of attributions. For example, both Integrated Gradients and TracIn admit inspiration from the Fundamental Theorem of Calculus.
Common approaches have also been utilized to evaluate the quality of all three types of attribution (Agarwal et al., 2022; Hammoudeh and Lowd, 2024; Rai et al., 2024). Counterfactual evaluation is a widely used approach, which compares the attribution scores with the actual impact of removing or modifying inputs. For example, evaluate by measuring the model’s performance drop when the inputs with high attribution scores are removed. Task-specific evaluation is another common approach to measure the attribution quality by running specific downstream tasks with the attribution results. In this type of evaluation, one can use FA to help identify feature changes that can flip model outputs, apply DA to detect mislabeled training examples, and run CA to identify important components that allows for model pruning. Moreover, human evaluation in many cases can be a gold standard by relying on domain experts or users to assess the quality and interpretability of attributions. More details on these evaluation approaches and their pros and cons are provided in Section F.3.
4 Leveraging the unified perspective to advance research
Feature, data, and component attribution have been studied largely as separate research areas, resulting in the parallel development of similar methods from different communities with distinct terminologies. Having demonstrated the unified view of the three types of attribution, we now discuss how this unified view can help advance both interpretability and broader AI research. We point out open challenges as well as potential solutions as future research directions.
4.1 Advancing research in interpretability
4.1.1 Overcoming shared challenges among all attribution types
Attribution methods face several common challenges that need to be addressed to enhance their practicality and reliability. Our unified perspective provides clearer insights into the underlying causes of these challenges and their potential solutions. We discuss these key challenges and propose potential solutions both below and in greater detail in appendix F.
Improving computational efficiency. Computational efficiency is a significant challenge for all three types of attribution methods, preventing them from being applied to large models (Adolfi et al., 2024; Bassan et al., 2024; Barceló et al., 2020). Our unified view reveals that this challenge is rooted in their shared techniques. For perturbation-based methods, the curse of dimensionality makes it intractable to comprehensively analyze high-dimensional inputs, large training datasets, and models with numerous components. Gradient-based methods offer more practical computational costs, except when sophisticated gradient computations are needed, such as aggregating gradients or computing second-order Hessian matrices. Linear approximation methods also face computational challenges when numerous data points and model evaluations are required to establish sufficient data for learning accurate linear models. Therefore, improving computational efficiency calls for more efficient sampling methods (e.g., Maximum Sample Reuse (Wang and Jia, 2023)) to avoid perturbing the entire input space in perturbation-based methods and to learn more accurate linear models for linear approximation methods. Also needed are more efficient gradient computations, especially for higher-order gradients like Hessians. Many works have sought to improve these aspects, with promising examples like Grosse et al. (2023) showing that gradient-based methods can be scaled to large language models (LLMs) with up to 52 billion parameters. However, there is still much room for improvement to make attribution methods more efficient. Also, a unified perspective may be developed for the corresponding computational problems (Adolfi et al., 2024; Bassan et al., 2024; Barceló et al., 2020).
Improving attribution consistency. Consistency challenges refer to the variability in attribution results across multiple runs of the same method with different random seeds, making it challenging to establish stable interpretations and evaluations (Jeyakumar et al., 2020). Our unified view reveals that this challenge is prevalent across all three attribution types due to multiple of randomness rooted in their common techniques, including sampling, learning processes with stochastic optimization, and non-trivial hyperparameters. We see the potential improvement for this challenge including producing more deterministic attribution methods that are less sensitive to random hyperparameters, provide provable guarantees of attribution stability or error bounds. While past works have made some progress in this direction, there were strong assumptions added to the setting, e.g., convexity (Koh and Liang, 2017). Thus, more efforts are needed to improve consistency in more realistic settings.
Towards fairer and more practical evaluation. Evaluation challenges arise in all three types of attribution due to multiple factors. Inconsistent results make it difficult to reliably compare methods and determine their relative accuracy. This challenge is particularly evident in gradient-based FA methods. As demonstrated by Adebayo et al. (2018), they can produce contradictory attribution results and sometimes perform no better than random baselines, making them difficult to evaluate fairly. For the shared evaluation metrics among all three types of attribution discussed in Section 3.4, counterfactual evaluation could provide more rigorous validation, but computational constraints often make this approach impractical. Task-specific evaluations offer easier alternatives but frequently lack generalizability across different contexts. Human evaluation, despite being considered the gold standard, faces scalability issues and potential biases. The diversity of evaluation metrics and their varying definitions of importance make the evaluation more challenging. An attribution result may perform well by one metric but poorly according to another. These challenges emphasize the pressing need for developing more reliable and generalizable attribution evaluation metrics.
Altogether, these are open challenges that call for future research to develop more efficient and reliable attribution methods as well as rigorous and principled frameworks to evaluate them. Given the shared techniques and common challenges among different attribution types, ideas to address these challenges in one attribution type will likely also be relevant to others.
4.1.2 Cross-attribution innovation
The connections among FA, DA, and CA discussed in the sections above naturally suggest promising research directions of leveraging insights from one type of attribution to develop methods for another. This can be directly identified as filling in the empty cells in Table 1. For example, while the Shapley value has been successfully applied across all three types of attribution, many other game-theoretic notions have only been used in FA and not for DA and CA yet. In addition, some advanced gradient techniques are common for FA and DA methods, but not for CA methods. The Hessian matrix, for example, has been used to obtain second-order information in Integrated Hessians for FA and extensively in all IF-related methods for DA, and can be explored for CA.
Moreover, seeing the theoretical connections among FA, DA, and CA enables us to draw inspiration from one area to advance our understanding of another as a whole. For example, Section § 3.1.1 discusses how diverse FA methods all perform LFA. This framework may also apply to DA and CA. We know that FA performs function approximation of the blackbox model’s predictions over the space of input features. One may hypothesize that DAs perform function approximation of the model’s weights over the space of training data points and that CAs perform function approximation of the model’s predictions over the space of model components. If so, function approximation may unify DA methods and unify CA methods as well. As in the case of unifying FA methods, such theoretical unification may also provide many benefits to DA and CA, including conceptual coherence, elucidation of method properties, simplicity for theoretical analyses, and principled practical recommendations.
Another research direction is to move towards more holistic analyses of model behavior. These attribution methods provide insight into model behavior through different lenses: input features, training data, and model components. Each type of attribution provides different and complementary information about model behavior. For example, for a given model prediction, FA may not suggest that the model is relying on sensitive features to make predictions, but CA may uncover a set of neurons that encode biased patterns. In this sense, focusing only on one type of attribution, i.e., studying only one part of the model, is insufficient to understand model behavior. Thus, future research may develop approaches to enable more comprehensive model understanding, such as understanding how to use different types of attribution methods together, the settings under which different attribution types may support or contradict one another, and the interactions between the three model parts (e.g., how patterns in the training data are encoded in model neurons).
4.2 Advancing research in broader AI
Attribution methods also hold immense potential to benefit broader AI research for other applications. Especially with a unified view integrating FA, DA, and CA, researchers cannot only gain deeper insights of model behavior but also edit and steer models towards desired goals and improve model compliance with regulatory standards.
Model editing (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022, inter alia) focuses on precisely modifying models without retraining. It enables researchers to correct model mistakes, analogous to fixing bugs in software. This approach is particularly valuable for LLMs, which encode vast information in their parameters and are prohibitively expensive to retrain. It can be viewed as a downstream task of attribution methods. Once attribution methods locate an issue, editing methods can be applied to the problematic parts. While editing aligns most closely with CA, other attribution types serve essential complementary functions. FA identifies spurious correlations requiring correction, and DA helps select proper data points to induce efficient editing. The unified attribution framework provides a holistic perspective that enables more efficient and accurate editing, especially when CA alone proves insufficient (Hase et al., 2024).
Model steering (Zou et al., 2023; Marks and Tegmark, 2023; Bürger et al., 2024, inter alia) differs from model editing by integrating a steering vector into the model’s inference process rather than modifying model parameters. While editing focuses on specific knowledge modifications, steering guides model behavior at a higher level, such as enhancing truthfulness and harmlessness in LLMs. Similar to model editing, a unified attribution framework can significantly enhance steering by better localizing target components to steer and generating more effective steering vectors through relevant features and training data. Recently, steering methods that rely on more than one type of attribution methods have emerged Marks and Tegmark (2023); Bürger et al. (2024). In Appendix G, we provide a more detailed discussion of how a unified attribution framework can enhance both model editing and steering and provide a illustrative example of steering for truthfulness of LLMs.
Model regulation (Oesterling et al., 2024, inter alia) is an emerging field examining the relationship between AI systems, policy, and societal outcomes. Regulation and policy frequently stress the need for transparency of AI systems as well as users’ right to an explanation. Attribution methods provide an avenue for practitioners to ensure that AI systems meet these legal and ethical requirements, by providing information about the overall AI system as well as specific input-output behavior. FA reveals input processing patterns, DA exposes training data influences, and CA illuminates architectural roles. This multi-faceted understanding enables more targeted and effective regulation. For example, when addressing biased behavior, FA can be used to identify discriminatory input patterns, DA to trace problematic training samples or copyright infringements, and CA to locate architectural elements needing adjustment. These complementary perspectives provide the comprehensive understanding needed to guide model regulation toward desired societal outcomes.
5 Alternative views
Several alternative views exist that (partially) oppose the unified perspective presented in this paper. We discuss their merits and connections to our work below.
One alternative view holds that different attributions represent distinct research areas that have benefited from independent evolution in the past, and although they share some ideas, they will continue to benefit from developing separately. FA has historically been synonymous with ”explainable AI,” DA has been part of ”data-centric AI” focusing on applications such as data selection, and CA has emerged as a branch of ”mechanistic interpretability” aimed at reverse-engineering algorithms in foundation models. This independent evolution has produced distinct motivations and best practices within each area (Guidotti et al., 2018; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024). While this view does not strongly oppose unification, attribution methods can indeed be more easily compared with other works from their respective fields when viewed separately. For example, FA can be compared with inherently interpretable models (Koh et al., 2020) for feature importance, DA can be compared with data coresets (Mirzasoleiman et al., 2020) for data selection, and CA can be compared with model pruning (Sun et al., 2023). Nevertheless, we demonstrate in the second part of our position that meaningful unification across attributions is possible and can inspire new research directions. We emphasize the unified perspective and its benefits throughout this paper.
There are also alternative views on the taxonomy of attribution methods. First, researchers have proposed other taxonomies within each type of attribution. For example, FA methods have been classified as either model-agnostic or model-specific (Arrieta et al., 2020). Although these categorizations aid method selection, they are not applicable across all attribution types. CA methods, for instance, are inherently model-specific since they analyze defined model components. Second, one might argue that our three-way categorization of FA, DA, and CA into perturbation-based, gradient-based, and linear-approximation methods does not apply to all attribution methods. While this is true – there are surrogate models for approximation beyond linear (Sokol and Flach, 2020) for FA (Section C.5), simulation-based (Chai et al., 2024) and reinforcement learning-based (Yoon et al., 2020) methods for DA (Section D.6), and sparse autoencoder-based methods for CA (Bricken et al., 2023) (Section E.3) – our categorization captures the majority of representative methods. In summary, our unification extends beyond attribution-specific categorizations and highlights broader, cross-attribution similarities. It is relatively comprehensive and uncovers connections that remain obscured when examining attributions in isolation. For example, while methods from all three attribution types use gradients, they compute them for different purposes with different objectives.
Recent work by Saphra and Wiegreffe (2024) offers seemingly another alternative view on attributions, highlighting that different communities assign distinct meanings to attributions, often guided by historical, cultural, and methodological differences. However, by recognizing that attributions carry community-specific definitions with epistemic commitments, rather than being strictly independent methodologies, this view is actually aligned with our unification position.
6 Conclusion
In this paper, we present a unified view of attribution methods from explainable AI (e.g., FA), data-centric AI (e.g., DA), and mechanistic interpretability (e.g., CA). We demonstrate that, while these attributions evolved largely independently to explain different aspects of AI systems, they differ mainly in perspective (i.e., examining input features, training data, or model components) and share core techniques (i.e., perturbations, gradients, and linear approximations). This unified view not only bridges the fragmented landscape of attribution methods but also inspires new directions in interpretability research through addressing common challenges and cross-attribution innovation and in broader AI research through model editing, steering, and regulation.
7 Acknowledgments
Usha Bhalla is supported by a Kempner Institute Graduate Research Fellowship. We thank Suraj Srinivas, Yuanzhou Chen, Karim Saraipour, and Gaotang Li for their feedback on the draft of this work.
Appendix
Appendix A Summary of notations
In Table 2, we summarize the notation used in this paper.
| Notation | Description |
| Training dataset | |
| / | Model trained on , parameters may be omitted |
| Internal model components , definition is method-specific | |
| / | Model input at test time for inference, superscript “test” may be omitted |
| The -th input feature out of features | |
| Attribution score of input feature for model output | |
| Attribution score of training data point for model output | |
| Attribution score of internal model component for model output | |
| Attribution function, which provides attribution scores for elements | |
| Loss function for training the model | |
| Loss function for learning the attribution function |
Appendix B Summary of methods
As we discussed in the main paper, methods for feature, data, and component attribution can be categorized based on their technique: perturbation-based methods, gradient-based methods, and linear approximation methods. Perturbation-based methods measure how a model’s output changes when modifying specific elements, whether they are input features, training data points, or model components. To capture interactions between multiple elements, all three types of attributions employ common mathematical tools, such as the Shapley value from game theory. Gradient-based methods analyze model behavior by leveraging gradients to provide insights into the model’s sensitivity to small input changes. Gradients bridge model behavior and the elements we wish to attribute without perturbations. Attributions are generated using different kinds of gradients: computing gradients of model outputs with respect to input features to quantify feature importance, calculating gradients of loss functions with respect to specific training data points to analyze data influence, or using gradients to approximate the effects of modifying model components. Linear approximation methods fit linear models to approximate complex model behaviors. The inputs to these linear models can be input features, training data points, or model components. In some cases, binary indicators replace the actual elements as inputs to simplify the approximation.
In Table 1, we summarize the attribution methods discussed in this paper, which we believe are the representative ones and align with the unified view we presented. Most of the empty cells in Table 1 (labeled as “–”) represent methods that may be promising but have not yet been explored in the literature, as discussed in § 4.1.2. These represent research ideas that have been verified in one attribution type but remain unexplored in others. The exception is mask learning for data attribution, which may be less promising because learning a high-dimensional mask of size jointly with the model would be infeasible when the model has not been trained.
Appendix C Detailed discussion of feature attribution methods
Feature attribution methods can be broadly classified into three categories. Perturbation-based methods attribute feature importance by observing changes in model output when input features are altered or removed. These methods provide intuitive results but can be computationally expensive for high-dimensional data. Gradient-based methods utilize the model’s gradients with respect to input features to attribute their importance. These methods are popular for differentiable models like neural networks, as they are often computationally efficient. Linear approximation methods construct interpretable linear models of input features that approximate the behavior of the original complex model in the vicinity of a specific input and compute attribution scores from the linear model coefficients. These methods offer a balance between interpretability and local accuracy. Each category of methods has its strengths and limitations, making the choice of method dependent on the specific model, data characteristics, and attribution requirements of the task at hand. We now extend the discussion of some methods mentioned in the main text in more detail and provide discussions of some additional methods.
C.1 Direct perturbation for feature attribution
RISE
[Petsiuk, 2018] is a direct perturbation method that addresses limitations of earlier methods like occlusion while expanding applicability to complex models. The method provides a systematic approach for assessing feature importance through efficient sampling and aggregation of perturbations. It operates by randomly masking different regions of the input image and measuring the model’s output to each masked version. The final saliency map is constructed by combining these random masks, with each mask weighted according to the model’s predicted probability on the corresponding masked input. This sampling-based approach allows RISE to efficiently estimate feature importance while capturing interactions between different image regions.
C.2 Game-theoretic feature attribution
The Shapley value, a solution concept from cooperative game theory introduced by Lloyd Shapley [Shapley, 1953], has gained particular prominence in feature attribution. For a data point with features , the Shapley value of feature for the model prediction is defined as:
where represents a subset of features excluding feature indexed by , and denotes the model’s prediction when only features in set are present. is the marginal contribution of feature to the subset for the model’s prediction. The formula computes the average marginal contribution of feature across all possible feature subsets. We simplify the attribution score notation by writing .
Shapley values possess several desirable properties that make them particularly suitable for feature attribution:
Efficiency: The attributions sum to the total prediction, i.e.,
Symmetry: Features that contribute equally receive equal attribution, i.e., if for all subsets , then
Linearity: For models and , for constants and .
Null player: Features that don’t affect the prediction receive zero attribution, i.e., if for all subsets , then
These properties offer theoretical guarantees for fair and consistent feature attribution, making Shapley values a principled approach to understanding model behavior. However, the exact computation requires evaluating feature combinations, leading to various approximation methods in practice.
Other game-theoretic concepts
Besides the Shapley value, other cooperative game-theoretic concepts are also applicable to feature attribution, offering different trade-offs between computational complexity and specific properties of the resulting attributions. The Shapley Taylor Interaction Index (STII) [Dhamdhere et al., 2019] is another concept that can be used for feature attribution, which is a generalization of the Shapley value that explicitly considers interactions between features. The Banzhaf Interaction Index (BII) [Patel et al., 2021] is particularly useful for considering joint feature interactions with simpler computation than the Shapley value. The core value [Yan and Procaccia, 2021], for instance, employs different axioms and emphasizes attribution stability. Additionally, the Myerson value [Chen et al., 2018b] and HN Value [Zhang et al., 2022] are valuable when prior knowledge about the feature structure is available. These alternative approaches provide researchers and practitioners with a range of tools to tailor their feature attribution methods to specific needs and constraints of their models and datasets.
Connection to linear approximation
The most common Shapley value-based attribution method, SHAP [Lundberg and Lee, 2017], is a perturbation-based method rooted in cooperative game theory. However, it can also be viewed through the lens of linear approximation methods, representing a unified approach of local linear attribution and classic Shapley value estimation. In the context of linear approximation, SHAP can be interpreted as fitting a linear model where features are players in a cooperative game, and the model output is the game’s payoff. The SHAP framework includes variants like Kernel SHAP, which uses a specially kernel for weighted local linear regression to estimate SHAP values, effectively approximating the model’s behavior in the feature space surrounding the instance being explained. This perspective on SHAP highlights its connection to linear approximation methods while retaining its game-theoretic foundations.
C.3 Perturbation mask learning for feature attribution
Mask learning methods offer several notable advantages in feature attribution. They provide attributions more efficiently, especially for high-dimensional inputs, which is particularly beneficial for complex models and large datasets. The continuous spectrum of importance scores generated by these techniques offers more nuanced insights than binary approaches, allowing for a finer-grained understanding of feature relevance. Furthermore, the learning process can implicitly capture complex feature interactions, providing a more comprehensive view of how features contribute to model decisions. Additionally, these methods can be tailored to specific model architectures and incorporate domain-specific constraints, enhancing their flexibility and applicability across various fields.
Masking model
[Dabkowski and Gal, 2017] pioneered the masking approach for image classification by introducing a masking model that generates pixel masks, aiming to identify a minimal set of features that sufficiently maintain the prediction of the original input. This approach employs a U-Net [Ronneberger et al., 2015] as the masking model, which first encodes the input image and then upsamples it to produce a mask. The masking model acts as the attribution function , where mask values represent feature attribution scores. While initial training is required, the masking model generates masks through a single forward pass at test time, which significantly improves runtime compared to earlier perturbation methods. For learning the mask model, the main challenge is balancing feature minimality and predictive power.
L2X
[Chen et al., 2018a] frames feature attribution as an optimization problem and learns a masking model to generate masks that maximize mutual information between input feature subsets and model output. This approach not only identifies important features but also captures their interdependencies, providing a comprehensive explanation of the model behavior. L2X is versatile and applicable to various domains beyond image classification. For instance, it has been successfully applied to sentiment classification tasks using datasets of movie reviews.
Gradient computation in mask learning vs. gradient-based methods
To avoid potential confusion, it is important to note that while mask learning methods may utilize gradient computation during the learning process, these gradients serve a different purpose than those in gradient-based attribution methods. In mask learning, gradients are used to learn a soft mask or an explainer model for generating masks. These gradients are not used to directly determine the feature attribution scores themselves. This distinction sets mask learning approaches apart from the gradient-based methods discussed in the following section.
C.4 Gradient-based feature attribution
Gradients for feature attribution are very different from those used in model training. For a model with parameters , gradients of the loss function are taken with respect to the parameters () to guide parameter updates during training. In feature attribution, gradients of the model’s output are taken with respect to input features () to quantify each feature’s contribution to the model’s output. Gradient-based methods have been widely adopted for feature attribution because of their computational efficiency. They typically require only a single forward and backward pass through the model to compute the gradients, and require no additional perturbation or linear model fitting, which makes them particularly suitable for real-time applications and large-scale datasets. On the other hand, gradient-based methods have two key limitations. They require access to model parameters and only work with differentiable models. Additionally, the gradient results can be nonrobust as we discussed in Section 4.1.1. Despite these challenges, gradient-based methods remain a fundamental tool in the feature attribution toolkit.
Gradient Input
[Shrikumar et al., 2017] improves over vanilla gradients. By multiplying the input features element-wise with their corresponding gradients, this method mitigates the “gradient saturation” problem where gradients can become very small even for important features. The element-wise multiplication also helps reduce visual diffusion in the attribution score visualization, resulting in sharper and more focused visualizations of important features.
Integrated Gradients
[Sundararajan et al., 2017] provides a theoretically grounded approach to feature attribution by accumulating gradients along a path from a baseline input to the actual input. This method satisfies important axioms including sensitivity (a change in input leads to a change in attribution) and implementation invariance (attributions are identical for functionally equivalent networks). The integration process captures the cumulative effect of each feature as it transitions from the baseline to its actual value, providing a more complete picture of feature importance than vanilla gradients of a single input.
Integrated Hessians
[Janizek et al., 2021] extends the integrated gradients method to analyze feature interactions, with the goal of understanding how features interact. This method treats the integrated gradient function as differentiable and quantifies interactions between two features using second-order information, the Hessian matrix. By computing these Hessian-based interactions along the same integration path used in integrated gradients, it provides a principled way to measure feature interdependencies and more comprehensive feature attributions.
Guided Backpropagation (GBP)
[Springenberg et al., 2014] modifies the standard backpropagation process of NNs to generate cleaner and more interpretable attribution results. When propagating gradients through ReLU units, GBP sets negative gradient entries to zero, effectively combining the signal from both the higher layer and the ReLU units. This modification helps eliminate artifacts and noise in the attribution results while preserving the positive contributions of features, resulting in sharper and more visually interpretable feature attributions.
Grad-CAM
[Selvaraju et al., 2016] is a widely used method for attribution and visualization of important regions in images, specifically for convolutional NNs. It computes the gradient of the target class score (logit) with respect to the feature maps in the last convolutional layer. These gradients are then used as weights to combine the feature maps, creating a coarse localization map that highlights important regions for predicting the target class. The resulting localization map is upsampled to match the input image size to create a saliency map, providing an interpretable visualization of features important for the model’s output.
Generalizing gradient-based feature attribution
Several additional methods share similar underlying principles with gradient-based approaches, although they do not directly compute gradients in their original formulation. For example, Layer-wise Relevance Propagation (LRP) [Bach et al., 2015] propagates predictions backwards through the network while preserving the total relevance at each layer. LRP provides a unique perspective on attribution by focusing on the relevance of individual neurons to the final prediction. Similarly, DeepLIFT [Shrikumar et al., 2017] operates by comparing each neuron’s activation to a reference activation and propagating the resulting differences to the input features. Interestingly, Ancona et al. [2017] demonstrated that for ReLU networks with zero baseline and no biases, both -LRP and DeepLIFT (rescale) methods are mathematically equivalent to the Input Gradient approach. For a more detailed analysis of these equivalences, we refer readers to Ancona et al. [2017].
C.5 Linear approximation for feature attribution
C-LIME
[Agarwal et al., 2021] is a variant of LIME specifically designed for continuous features that generates local explanations by sampling inputs in the neighborhood of a given point. It differs from LIME in several aspects: it uses a constant distance metric and Gaussian sampling centered at the input point rather than uniform random sampling, making perturbations naturally closer to the input without requiring explicit weighting. C-LIME also restricts itself to linear models for continuous features, unlike LIME’s more general model class, and excludes regularization by setting the regularizer term to zero. For simplicity, C-LIME focuses on feature weights while ignoring the intercept terms.
Generalizing linear approximations for feature attribution
For both LIME, C-LIME, and other similar linear approximation methods, the assumption that the model’s behavior can be reasonably approximated by a linear function in the local neighborhood is crucial. In this context, while the linear model serves as a proxy, it can be replaced by a more complex yet interpretable model that is still capable of providing attribution results. For instance, LIMETree uses tree models [Sokol and Flach, 2020]. We refer readers to Sokol and Flach [2020] for a detailed discussion of this approach.
C.6 Unifying feature attribution methods through local function approximation
Under the local function approximation framework, the model is approximated by an interpretable model class around the point of interest over a local neighborhood distribution using a loss function . The approximation is given by
Han et al. [2022] show that at least eight feature attribution methods (Occlusion, KernelSHAP, Vanilla Gradients, Gradients Input, Integrated Gradients, SmoothGrad, LIME, and C-LIME) are all instances of this framework. These methods all use the linear model class to approximate , but do so over different local neighborhoods using different loss functions as in Table 3.
Under this setup, ’s model weights are equivalent to the explanation obtained using each method’s original algorithm. Also, note that for the local function approximation framework, there are requirements on the loss function: a valid loss is one such that .
In addition, while the local function approximation framework may seem similar to LIME, it differs from LIME by 1) requiring that and share in the same input and output domain, 2) imposing the condition on the loss function discussed above, and 3) following the standard machine learning methodology to avoid overfitting and to tune hyperparameters. A more detailed discussion can be found in Section 3 of Han et al. [2022].
| Techniques | Attribution Methods | Local Neighborhood around | Loss Function |
| Perturbations | Occlusion | Squared Error | |
| KernelSHAP | Squared Error | ||
| Gradients | Vanilla Gradients | Gradient Matching | |
| Integrated Gradients | Gradient Matching | ||
| Gradients Input | Gradient Matching | ||
| SmoothGrad | Gradient Matching | ||
| Linear Approximations | LIME | Squared Error | |
| C-LIME | Squared Error |
Appendix D Detailed discussion of data attribution methods
Unlike feature attribution methods, data attribution methods examine the training phase and quantify the training data’s influence on the model’s output. However, data attribution methods can be categorized into the same three categories. Perturbation-based methods assess training data importance by observing changes in model behavior when training samples are removed or modified. These methods provide accurate results but can be computationally expensive as they require retraining the model multiple times. Complete retraining-based methods like LOO are often used as a ground truth for evaluating other data attribution methods. Gradient-based methods utilize the model’s gradients evaluated at the training data points and the test data point to quantify the influence of the training data points on the test data point. These methods avoid the computational cost of retraining the model but may face the challenges like the non-convexity of the loss landscape or the difficulty in computing the Hessian matrix efficiently. Linear approximation methods construct interpretable models that approximate how training data affects model behavior. The linear model operates on the entire training dataset, which can be suprisingly accurate but also heavy to train. Methods from different categories have their own strengths, limitations, and use cases. We now extend the discussion of some prominent methods mentioned in the main text and discuss some additional methods.
D.1 Leave-One-Out data attribution
Leave-One-Out is a prominent example of perturbation-based data attribution, but also a natural idea that existed for a long time in statistics. For example, it has been used as a resampling technique (e.g., jackknife resampling [Tukey, 1958]) to estimate the bias and variance of a statistic of interest (such as a regression coefficient). LOO has been used to detect influential data points for linear regression [Cook and Weisberg, 1982], for example, through Cook’s distance [Cook, 1977]. Until recently, LOO has been applied to modern AI models to attribute model performance to individual training data points [Jia et al., 2021]. It provides valuable counterfactual insights with its main limitation being computational cost, as it requires retraining the model for each data point. Many newer attribution methods, like the gradient-based methods can be viewed as efficient approximations of LOO.
D.2 Game-theoretic data attribution
The primary limitation of game-theoretic methods is their prohibitive computational cost for large datasets, as they require numerous model retrainings over the powerset of the training data. To address this challenge and further improve method robustness, researchers have proposed various methods.
Truncated Monte Carlo (TMC) Shapley
[Ghorbani and Zou, 2019] approximates the Shapley value by adopting an equivalent definition of the Shapley value in terms of aggregating over data permutations instead of data subsets [Shapley, 1953]. The method works by truncating the number of permutations sampled and the number of data points considered in each permutation, and align that with model training. For each sampled permutation, it computes the marginal contribution of each data point by evaluating model performance with and without that point. The gradient information from these evaluations is then used to as an estimate of each point’s marginal contribution. TMC Shapley significantly reduces computational cost while maintaining reasonable approximation accuracy of the exact Shapley values.
KNN Shapley
[Jia et al., 2019] introduces an efficient approximation for Shapley values data attribution by using K-Nearest Neighbors (KNN) as a surrogate model instead of retraining the full model. For each test point, it first identifies its K nearest neighbors in the training set. Then, it computes Shapley values only considering these neighbors’ contributions to the KNN prediction, rather than the original model’s prediction. This localized computation dramatically reduces complexity from exponential to polynomial in the number of neighbors K. The method maintains good attribution quality since nearby training points typically have the biggest influence on a test point. The KNN approximation aligns better with the goal of estimating the value of data from the data vendor’s perspective, and thus was named data valuation in the original paper.
Beta Shapley
[Kwon and Zou, 2022] extends the standard Data Shapley framework by introducing a beta distribution to weight different subset sizes differently. The original Shapley value weights subsets according to their sizes. The parameter controls how much emphasis is placed on smaller versus larger subsets when computing marginal contributions. This generalization relaxes the efficiency axiom of classical Shapley values, which requires attributions to sum to the total model value. By allowing this flexibility, Beta Shapley can better handle noisy or corrupted training data by reducing their influence on the attribution scores. The method provides theoretical analysis showing how different values affect properties like noise robustness and estimation variance.
Data Banzhaf
[Wang and Jia, 2023] adapts the Banzhaf value from cooperative game theory as an alternative to Shapley values for data attribution. The Banzhaf value considers the average marginal contribution of training point across all possible data subsets like the Shapley value, but weights these contributions differently. This weighting scheme leads to the largest possible safety margin, making the attribution more robust to data perturbations and noise. The method provides theoretical guarantees on this robustness and demonstrates empirically that it can better identify mislabeled or adversarial training examples compared to Data Shapley.
D.3 Influence function and its variants
Influence functions provide a way to estimate how model parameters would change if we reweight or remove a training point, without having to retrain the model. Given a model with parameters trained by minimizing the empirical risk over the training, the IF approximates the change in parameters when upweighting a training point by :
Under the assumption that the loss function is twice-differentiable and strictly convex, a first-order Taylor expansion around the final optimal model parameters gives:
where is the Hessian and is by assumption positive definite. The influence of training point on the loss at test point is the effect of this infinitesimal -upweighting on test point’s risk:
The negative will be the data attribution score on and it provides an efficient approximation to LOO retraining [Koh and Liang, 2017].
While effective in certain scenarios and computationally more feasible than retraining-based methods, IF faces several challenges. First, it assumes convexity and double-differentiability, which are often not satisfied in deep learning scenarios. Second, it involves Hessian matrix computation, which can be computationally expensive for large models. Also, the potential non-positive definiteness of the Hessian matrix in certain cases can lead to inaccuracies, often necessitating the introduction of dampening factors that may affect the precision of influence estimates. To address these limitations, many methods have been proposed to enhance the efficiency and applicability of IF.
FastIF
[Guo et al., 2021] introduces several key optimizations to make IF more computationally tractable. First, it uses KNN to reduce the search space from the entire training set to a smaller subset of promising candidates that are likely to be influential. Second, it develops a fast estimation technique for the inverse Hessian-vector product that avoids computing and storing the full Hessian matrix and its inverse. Third, it implements parallelization strategies to asynchronously compute Hessian-vector products across multiple processors. These optimizations together enable FastIF to scale to much larger datasets while maintaining attribution quality comparable to the original IF.
Arnoldi IF
[Schioppa et al., 2022] employs Arnoldi’s iterative algorithm to efficiently identify the dominant eigenvalues and eigenvectors of the Hessian matrix. These dominant components serve as the basis for projecting all gradient vectors into a lower-dimensional subspace. Compute IF in this subspace substantially reduces the computational complexity. The method can be flexible by selecting an appropriate number of eigenvalues to retain. Empirical results demonstrate that this approach can achieve comparable attribution quality to full IF while significantly reducing both memory requirements and computation time.
EK-FAC
[Grosse et al., 2023] leverages the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) parameterization to efficiently approximate the Hessian matrix. This parameterization exploits the natural block structure present in NNs to decompose the Hessian into more manageable components, which significantly reduces the computational complexity of Hessian-vector products. By leveraging these techniques, IF can be effectively scaled to large transformer models with hundreds of millions of parameters, which are orders of magnitude more complex than the simpler NNs originally considered by Koh and Liang [2017]. Theoretical guarantees for the approximation quality and demonstrations of empirical success on foundation models were shown.
RelateIF
[Barshan et al., 2020] addresses another limitation of IF other than their computational cost. Standard IF methods often highlight outliers or mislabeled data points as most influential, which may not always align with intuitive notions of influence. RelateIF introduces a novel approach that distinguishes between global and local influence by examining how training data affect specific predictions relative to their overall impact on the model. This relative influence measure helps identify training data points that have significant local influence on particular test predictions while accounting for their broader effects on the model. RelateIF better captures intuitive notions of influence while being more robust to outliers in the training data.
D.4 Tracing training path for data attribution
Tracing training path for data attribution provides valuable insights of training dynamics while avoiding limitations of LOO and IF. These methods that trace training dynamics that provide more accurate attribution results but they are all more computationally expensive than those considering only final model parameters like IF.
TracIn
[Pruthi et al., 2020] traces the data influence throughout the entire training process. The method attributes influence by computing dot products between training and test data gradients at each training step from the initial model parameters to the final model parameters at the end of training, accumulating these to capture a training point’s total influence across the training path. This path tracing approach provides valuable insights into training dynamics while avoiding limitations of LOO and IF, such as assigning identical attribution scores to duplicate training data points. TracIn also offers greater flexibility than IF by eliminating the convexity assumption and Hessian matrix computations. On the other hand, its tracing requires storing intermediate model checkpoints during training, increasing both memory usage and computational costs.
SGD-Influence
[Hara et al., 2019] traces the training path by approximating the training process with a series of unrolled steps to estimate data influence. The method estimates LOO influence by unrolling gradient descent using empirical risk Hessians, under the assumption that both the model and loss function are convex and the optimization algorithm is Stochastic Gradient Descent (SGD). SGD-Influence primarily applies unrolling to quantify the Cook’s distance [Cook, 1977] between model parameters with and without a specific training point. To better align with attribution estimation, a surrogate linear influence estimator is used to incrementally update throughout the unrolling process. However, this approach requires unrolling the full training path for each test instance individually, which has significant computational complexity.
SOURCE
[Bae et al., 2024] extends training path tracing to better capture the training dynamics and reduce the computational cost. It bridges the gap between gradient-based approaches like IF and unrolling-based methods like SGD-Influence. While IF is computationally efficient, it struggles with underspecification of the training dynamics. Unrolling-based methods address these limitations but face scalability challenges. SOURCE combines the benefits of both approaches by using an IF-like formula to compute approximate unrolling. This makes SOURCE both computationally efficient and suitable for scenarios where IF struggles, such as non-converged models and multi-stage pipelines. Empirically, SOURCE demonstrates superior performance in counterfactual prediction compared to existing data attribution methods.
D.5 Linear approximation and Datamodel
Datamodel
[Ilyas et al., 2022] applies linear approximation to DA, similar to LIME in FA. It constructs a linear model with coefficients and vectors as inputs, where each input represents a subset of training data. is learned to map any counterfactual subset of training data to output , where is trained on this subset with the given model architecture and training algorithm. The coefficients of thus represent the attribution scores of the training data points. The method’s counterfactual nature enables evaluation of other attribution methods via the Linear Datamodeling Score (LDS), which compares their attribution score rankings to Datamodel’s ranking. While Datamodel can effectively capture model behavior, constructing this large linear model requires extensive counterfactual data obtained by training model on various subsets, making it computationally intensive.
TRAK
[Park et al., 2023] addresses the computational challenges of Datamodel by estimating Datamodels in a transformed space where the learning problem becomes convex and can be approximated efficiently. It further improves efficiency through random projection of model parameters and ensemble attribution results of multiple trained models. Though the ensemble approach still requires some model retraining on different subsets, it achieves high estimation accuracy with significantly fewer retraining iterations than Datamodel. Furthermore, both approaches can be viewed as perturbation-based methods, similar to LIME, as they systematically vary training data to construct linear models.
D.6 Data attribution methods using other techniques
There are data attribution methods that use other techniques beyond the three core techniques we discussed in the main paper to estimate the influence of each training data point on the model’s output.
GPTfluence
[Chai et al., 2024] is a simulation-based method that estimates how individual training examples influence the performance of GPT models. GPTfluence uses a featurized simulation-based approach to model the training trajectory. It begins by collecting training dynamics across multiple runs, including sequences of training batches and performance metrics like loss, BLEU, and ROUGE scores. GPTfluence then trains a simulator that represents both training and test examples using frozen pre-trained embeddings (e.g., from BERT), and learns multiplicative and additive influence factors through inner products between these embeddings. The simulator models performance at each step using an n-th order Markov process that conditions on previous steps, allowing it to predict the full evolution of test metrics over training.
Data Valuation using Reinforcement Learning (DVRL)
[Yoon et al., 2020] uses reinforcement learning to estimate the influence of each training data point on the model’s output. DVRL jointly trains a data value estimator and a task predictor model. At each iteration, the data value estimator assigns selection probabilities to training examples, determining which examples are used to update the predictor model. The quality of these selections is evaluated using a small validation set, and a RL signal based on the validation performance is used to update the data value estimator. This allows DVRL to dynamically learn which samples are most useful for improving the model’s performance.
Appendix E Detailed discussion of component attribution methods
Unlike feature and data attribution methods, component attribution methods analyze the internal mechanisms of models by attributing model behavior to specific architectural components like neurons, layers, or attention heads. These methods similarly fall into three categories. Perturbation-based methods assess component importance by observing changes in model behavior when specific components are modified, resulting in various forms of causal mediation analysis. Gradient-based methods utilize gradients with respect to component activations to approximate the component importance in causal mediation analysis. Linear approximation methods construct linear models that directly approximate how components affect model behavior. Methods from different categories have their own strengths and limitations. We now extend the discussion of some prominent methods mentioned in the main paper and provide discussions of some additional methods.
E.1 Perturbation-based component attribution
Various types of ablations in causal mediation analysis
Causal mediation analysis is frequently referred to as activation patching, wherein activations of the specific component from the clean run are patched into the corrupted run to ascertain if those activations are sufficient and necessary to retrieve the desired output. Activation perturbations can consist of zero ablations [Olsson et al., 2022, Geva et al., 2023], mean ablations [Wang et al., 2022], smoothed Gaussian noising Meng et al. [2022], interchange interventions [Geiger et al., 2021], learned ablations [Li and Janson, 2024]. In all cases, the dataset used to generate the activations must be chosen to elicit the desired model behavior, with a matching metric that measures the success of the behavior.
Automated Circuit Discovery (ACDC)
Similar to subnetwork pruning, ACDC [Conmy et al., 2023] is tries to find a subnetwork that is far sparser than the original graph and recovers good performance on the task. This is done by iterating through the computational graph of the model from outputs to inputs and attempting to remove as many edges between nodes as possible without reducing the model’s performance. In this case, performance is measured as the KL-divergence between the full model and the subgraph’s predictions. Furthermore, masked or ablated edges are replaced with activations from a corrupted run or counterfactual input prompt, rather than zero-ablated as is done in subnetwork probing.
E.2 Gradient-based component attribution
Edge Attribution Patching (EAP)
[Syed et al., 2023] combines ACDC and Attribution Patching to create EAP, which generates attribution scores for the importance of all edges in the computational graph through normal attribution patching and then sorts those scores to keep only the top edges in a circuit, thus yielding the circuit corresponding to the task.
E.3 Generalizing the definitions of components
While initial works explored this form of causal mediation analysis where each neuron was an individual component [Bau et al., 2020, Vig et al., 2020], recent work has moved towards other mediators due to the computational intractability of considering individual neurons in larger models and due to hypotheses of entanglement and polysemanticity of neurons in foundation models. Furthermore, recent work has argued that specific mediators are only reasonable for certain behaviors [Mueller et al., 2024] and have also explored the feasibility of patching activations both within and between models to increase expressivity [Ghandeharioun et al., 2024].
Sparse Autoencoders (SAEs)
[Bricken et al., 2023, Cunningham et al., 2023] are trained to reconstruct model activations under sparsity constraints. Through learning sparse, overcomplete representations of model activations, SAEs effectively decompose complex, entangled features into more interpretable components. The enforced sparsity ensures that each SAE feature captures a distinct and meaningful aspect of the model’s behavior, making them useful for model understanding. Recent research has demonstrated that SAEs can successfully extract interpretable components, but since SAEs focus on learning new components rather than attributing to existing ones in the original model, we do not consider them as strictly component attribution methods in this paper. They can rather serve as a technique for discovering interpretable features that can subsequently be used for attribution. As the next paragraph shows, SAEs can be used for component attribution by first discovering interpretable components and then using them for attribution.
Sparse Feature Circuits
[Marks et al., 2024] Sparse feature circuits build upon the gradient-based attribution method attribution patching to determine the linear directions relevant to the task or behavior of interest. This method leverages sparse autoencoders to find directions in the models’s latent space that correspond to human-interpretable features. They then employ linear approximations similar to attribution patching, using either input gradients or integrated gradients, to efficiently identify which of the learned sparse autoencoder features are most relevant to the model behaviors, as well as connections between these features.
Appendix F Overcoming attribution challenges through a unified lens
F.1 Improving computational efficiency
Computation challenges present substantial barriers that often prevent attribution methods from being applied to large-scale AI models, such as the foundation models with billions of parameters. For perturbation-based methods, the curse of dimensionality makes a comprehensive analysis intractable when the number of required perturbations is large. For example, the full power set perturbation. This holds for all three types of attributions including high-dimensional inputs, large training datasets, and models with numerous components. Game-theoretic methods face particular difficulties, as exact computation of the Shapley value is often prohibitively expensive and requires approximation techniques like Monte Carlo sampling. The computational burden is also severer for data attribution methods, which require model retraining for each perturbation. Gradient-based methods are more practical for large-scale models. However, gradients essentially only provide first-order approximations of model behavior, which are inadequate to capture complex model behaviors and more sophisticated gradient formulations are needed for better attribution results. For instance, TracIn require aggregating gradients across multiple stages, while IF demand computation of second-order Hessian matrices, leading to increased computational overhead. Linear approximation methods also face computational hurdles in achieving high-quality approximations. Model behavior can be complex, requiring numerous data points and model evaluations to establish sufficient data for learning accurate linear models. Furthermore, for all three types, most attribution methods must compute results separately for each new test data point, creating additional computational strain when attribution analysis is needed for large datasets. Therefore, improving computational efficiency calls for better sampling methods to avoid perturbing the entire input space in perturbation-based methods and to learn more accurate linear models for linear approximation methods. Also needed are more efficient gradient computations, especially for higher-order gradients like Hessians. Many works have sought to improve these aspects, with promising examples like Grosse et al. [2023] showing that gradient-based methods can be scaled to large language models (LLMs) with up to 52 billion parameters. However, there is still much room for improvement to make attribution methods more efficient.
F.2 Improving attribution consistency
The consistency problem in attribution methods is a significant concern. This challenge is also prevalent across due to variability introduced in sampling, learning processes with stochastic optimization, and also non-trivial hyperparameters. When attribution involves sampling, such as the Monte Carlo sampling in some perturbation-based methods to avoid the full power set perturbation, the inherent randomness leads to varying attribution results. Besides, when attribution involves learning processes with stochastic optimization, as seen in mask-learning perturbation and linear approximation methods, different learning outcomes yield inconsistent attribution results. Many attribution methods rely on hyperparameters that can lead to different attribution outcomes. These include sampling parameters, such as the number of samples used for computing Shapley values. They also include optimization hyperparameters for various learning approaches, such as learning rates and number of steps in linear approximation methods. Additionally, approximation hyperparameters are needed for quantities that are computationally challenging to calculate directly, such as dampening factors for inverse Hessian-vector products. Further variability is introduced through fundamental design choices, such as the selection of perturbation type in perturbation-based methods, where options include mean perturbation, zero perturbation, and random perturbation. While these different approaches should theoretically produce similar results based on their underlying principles, in practice they often yield notably different attributions. While some gradient-based methods can produce consistent results in a single run when they do not involve sampling or approximations for computationally intensive operations, the consistency among different gradient-based methods varies considerably, which is also a problem for all three techniques and leads to the following evaluation challenges. We see the potential improvement for this challenge including producing more deterministic attribution methods that are less sensitive to random hyperparameters, provide provable guarantees of attribution stability or error bounds. While past works have made some progress in this direction, there were strong assumptions added to the setting, e.g., convexity [Koh and Liang, 2017]. Thus, more efforts are needed to improve consistency in more realistic settings.
F.3 Towards fairer and more practical evaluation
Evaluating attribution methods presents significant challenges due to the lack of ground truth and the inherent complexity of modern AI systems. These challenges stem partially from the inconsistency problem, the computational cost and generalizability of some evaluation metrics, and the lack of universal definitions of importance and ground truth. These common evaluation approaches and their limitations are summarized below.
Counterfactual evaluation
is a widely used approach that assesses attribution methods by comparing their scores with the actual impact of removing or modifying elements. Common metrics include fidelity, which evaluates sufficiency by retaining only elements with high attribution scores while removing those with low scores. Conversely, inverse fidelity measures necessity by removing elements with high attribution scores while retaining those with low scores. LOO attribution represents a special case of inverse fidelity. For data attribution specifically, more sophisticated metrics like LDS for data attribution compare attribution rankings with the actual impact of removing training data points, with LDS being a sophisticated case of fidelity. An important implicit metric in counterfactual evaluation is sparsity or minimality, which measures how few elements are needed to achieve high fidelity. Greater sparsity is desirable as it indicates that fewer elements are required for explanation. While counterfactual evaluation provides concrete validation, it faces two major challenges: The computational cost of generating counterfactuals, particularly for data attribution, can be prohibitive. Additionally, the complex interactions between elements may not be fully captured by individual counterfactual evaluations.
Task-specific evaluation
assesses the practical utility of attribution methods in downstream tasks. For instance, feature attribution can help identify feature changes that can flip model outputs, while data attribution scores can detect mislabeled training examples, and component attribution scores can help identify the most important components that allows for model pruning. Attribution methods can be compared based on the performance on these specific tasks. While this approach provides practical validation, its findings may not generalize effectively across different tasks or domains.
Human evaluation
relies on domain experts or users to assess the quality and interpretability of attributions. This approach is especially valuable for validating whether attributions align with human understanding and domain expertise. For example, for feature attributions, the attribution results can be considered if they generate clearer visual saliency maps that align with human intuition. While human evaluation provides valuable real-world validation, it can be both subjective and resource-intensive and can only be treated as the gold standard in certain cases.
The development of more robust and comprehensive evaluation frameworks remains a crucial research direction for advancing all attribution methods.
Appendix G Advancing attribution benefits broader AI research
Model editing and steering
A unifying framework offers a principled foundation and enables coordinated use of different attribution types—for example, combining CA’s localization with DA’s data selection under consistent semantics—leading to more effective and generalizable interventions. While recent model editing and model steering approaches have started to selectively target features, data, or components for better results, these methods are often heuristic. A unified attribution perspective can make these decisions more systematic and effective.
We use model steering for truthfulness as an illustrative case. In the foundational RepE work Zou et al. [2023], steering is performed via Low-Rank Representation Adaptation (LoRRA) trained on a set of contrastive input pairs. In RepE, 1) The contrastive dataset is constructed using random samples from QA datasets, 2) the layer to steer is treated as a hyperparameter, and 3) the target token position to steer is implicitly the final token. We argue that DA, CA, and FA can each contribute to a more principled approach to these three steps respectively, through selecting contrastive examples, localizing the appropriate layer, and identifying the right token position. While the original choices in RepE are not necessarily flawed (e.g., steering at the last token might give the best results), a unified attribution perspective offers a more principled framework with room for improvement.
Subsequent work supports this idea. For instance, Marks and Tegmark [2023] applies causal tracing to identify the most intuitive layer and token position for steering—an application of CA/FA ideas. Furthermore, Bürger et al. [2024] shows that the steering direction identified in Marks and Tegmark [2023] is not universal: models trained on affirmative vs. negated statements yield different truthfulness directions. Recognizing this limitation essentially involves a manual form of DA.
While the improvements in Marks and Tegmark [2023] and Bürger et al. [2024] are already promising, we believe that a unified attribution view can further enhance them. For example, the main-mass method used in Marks and Tegmark [2023] and Bürger et al. [2024] identifies steering directions by cluster centers, requiring application to a whole dataset. In contrast, principled data selection via DA could find effective steering directions with fewer samples. Importantly, these selected samples will be most effective when they reflect the truthfulness through critical components and features—highlighting the need for a holistic attribution approach.